OpenStudy (anonymous):

Help please!!!! 3x + 4y = 24 in a x and y chart fill in the blanks

OpenStudy (anonymous):



OpenStudy (anonymous):

x=2 , 3(2)+4y=24 6 + 4y=24 4y= 24-6 y=18/4

OpenStudy (anonymous):

what

OpenStudy (khaakenajaf):

3x + 4y = 24 4y=24-3x y=(24-3x)/4 now put values of x one by one for 1st x=2 y=(24-4(2))/4 y=(24-8)/4 y=16/4=4 now do the rest